Operations Report – December 5th

SOL: 8

Name of person filing report: Ashley Kowalski

Non-nominal systems: Toilet Issues

Notes on non-nominal systems:

The Health and Safety Officer, Barbara Braun, performed troubleshooting of the toilet. The holding tank for the toilet has been filling more quickly than expected and smelling bad for longer than expected after emptying. Upon advice of Tier 1 support, the HSO first emptied the holding tank into the septic tank, then flushed the holding tank with approximately two gallons of water. Then the HSO examined the inside of the tank with a headlamp and used a stick to estimate the level of material remaining in the tank after flushing. There were approximately two inches of watery sludge in the bottom of the holding tank even with the valve to the septic tank open. Upon the advice of Tier 1 support, the HSO then closed the septic valve, filled the tank full of fresh water until the water came up to the level of the bowl, and emptied the holding tank again by opening the septic valve. The results were essentially the same; approximately 1-2 inches of watery sludge still remained in the bottom of the holding tank even after the full water flush. Tier 1 support then advised the HSO to fill the tank with approximately five gallons of fresh water and add triple the usual amount of digestive. HSO accomplished this. Note that this activity used up one of the remaining jugs of digestive; we still have a full jug available. Tier 1 support will investigate further and consider servicing the holding tank in a break between rotations.


Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: Yes
Hours: 215.3 (Start), 215.6 (End)
Beginning charge: 100%
Ending charge: 70%
Currently charging: No.

Perseverance rover used: Yes
Hours: 250.5 (Start), 250.7 (End)
Beginning charge: 100%
Ending charge: 56%
Currently charging: No.

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 300 used (~250 gallons remaining – This is my best guess, but I am requesting Mission Support to let us know a more accurate value, if possible.)
Water (static tank): The anomalous toilet situation today caused the crew to use additional water resources. Tier 1 support provided an additional 120 gallons to make up for this loss.
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes. Special instructions that were provided to the crew by Mission Support for the anomalous toilet situation were followed. Please reference the “Non-nominal systems” section for details.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer performed the additional steps that were suggested to her by Mr. Scott Davis from the EVA suit support team. The Crew Engineer fully charged suits 3, 4, and 6 so that all battery charges were >12 V. She then ran each of those suits for about an hour. Suits 3 and 4 maintained a voltage >12 V. However, Suit 6 dropped to about 10.41 V. She left Suit 6 recharging for several hours, yet the battery did not charge up again to >12 V. In fact, after several hours of charging, the suit only charged to ~10.47 V. The suit seemed to still be running normally at that voltage. Thus, the suit seems to still be functional, but there might indeed be a slight battery issue with Suit 6. Please advise.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~39 gallons (~5 gallons today)
Heater: On
Supplemental light: On
Harvest: 17 grams of microgreens and 11 grams of volunteer plants

Summary of ScienceDome operations: Microscope analysis of the mirror samples were performed using The Aerospace Corporation microscope in the ScienceDome today.
Dual split: The heat is on in the ScienceDome and set to 65 F. Please note that as the Crew Engineer was performing mirror sample inspections today in the ScienceDome, it was quite cold still.

Summary of RAM operations: The multimeter from the RAM was used to check the voltages of the two-piece EVA suits during today’s inspections. Additionally, a large wooden paint stick (to assist with the toilet issues), the pulse oximeter, rubbing alcohol, and several sharpies were retrieved from the RAM.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: Heat furnace filter has not yet been installed. Please advise if this is something that the Crew Engineer needs to do.

Copyright © The Mars Society. All rights reserved. | Main Site